OpenStudy (chlobug):

Hello, everyone! I need really appreciate some help with a chemistry assignment. The homework question is below; no rush! Calculate the mass of ammonium sulfide (NH4)2S in 3.00 L of a 0.020 M solution.

OpenStudy (aaronq):

Find the moles then the mass \(\sf Molarity=\dfrac{moles}{L}\) then \(\sf moles=\dfrac{mass}{Molar~mass}\)

OpenStudy (chlobug):

Thanks, Aaron! I had a bunch of equation written down, but they were confusing me. I'll try that! (: Thanks again!

OpenStudy (aaronq):

no problem!

OpenStudy (chlobug):

This is what I did so far: \[M = 0.020\div3 = .0066\] Then: \[moles = .0066\div68.154 = 9.7\] Is my answer correct, along with my significant figures @aaronq? Thanks (:

OpenStudy (aaronq):

the first step is wrong, you solved for molarity instead of moles.

OpenStudy (aaronq):

rearranged, the formula is \(\sf moles=Molarity*L\)

OpenStudy (chlobug):

Oh, okay. What should the first equation be with numbers plugged in? I don't even have the moles yet, do I? I think I mistook the M (molarity) for moles. In order to find the moles, do I use this equation, @aaronq ? \[grams = mol/gram\] So, converting liters to grams then finding molar mass? \[3000 g = 1 mol/ 68.1 g (NH4)2S\]

OpenStudy (aaronq):

nope. you use the equation i posted originally. You were given molarity and volume (in L \(\sf Molarity =\dfrac{moles}{L}\rightarrow moles=Molarity*L=0.020 ~M* 3 L=0.060 ~moles\) then \(\sf moles=\dfrac{mass}{Molar~mass}\rightarrow mass=Molar~Mass*moles=0.060 ~moles*68.1 ~g/mol\)

OpenStudy (aaronq):

btw, "grams=mol/gram" is not correct. how can that be equal?

OpenStudy (chlobug):

Thank you! I got 4.08, but unfortunately forgot to round it up to 4.09. Thank you so much, @aaronq! I'm really sorry for all of my questions.

OpenStudy (aaronq):

No worries, that's why were here. I'm glad you got it figured out though!